# Exact Diagonalization, Matrix Product States and Conformal Perturbation Theory Study of a 3D Ising Fuzzy Sphere Model

### v1

@@ Abstract
Numerical studies of phase transitions in statistical and quantum lattice models provide crucial insights into the corresponding Conformal Field Theories (CFTs). In higher dimensions, comparing finite-volume numerical results to infinite-volume CFT data is facilitated by choosing the sphere $S^{d-1}$ as the spatial manifold. Recently, the fuzzy sphere regulator Zhu2023 has enabled such studies with exact rotational invariance, yielding impressive agreement with known 3D Ising CFT predictions, as well as new results. However, systematic improvements and a deeper understanding of finite-size corrections remain essential. In this work, we revisit the fuzzy sphere regulator, focusing on the original Ising model, with two main goals. First, we assess the robustness of this approach using Conformal Perturbation Theory (CPT), to which we provide a detailed guidebook. We demonstrate how CPT provides a unified framework for determining the critical point, the speed of light, and residual deviations from CFT predictions. Applying this framework, we study finite-size corrections and clarify the role of tuning the model in minimizing these effects. Second, we develop a novel method for extracting Operator Product Expansion (OPE) coefficients from fuzzy sphere data. This method leverages the sensitivity of energy levels to detuning from criticality, providing new insights into level mixing and avoided crossings in finite systems. Our work also includes validation of CPT in a 1+1D Ising model away from the integrable limit.
@@ 1 Introduction
Numerical studies of continuous phase transitions in statistical and quantum lattice models are a valuable source of information about the corresponding Conformal Field Theories. In numerical studies done in finite volume, extracting infinite-volume CFT data may be more or less difficult depending on the choice of the spatial manifold. For 1+1D models, studies of quantum Hamiltonians and transfer matrices on a circle $S^{1}$ allow for easy comparison to CFT thanks to the radial quantization Cardy:1984rp ; BCN ; Affleck ; finite-size-scaling . For higher-dimensional models, analogous simplicity would arise for the spatial manifold being the sphere $S^{d-1}$ Cardy_1985 . This was studied for 3D lattice models Brower:2012mn ; Brower:2012vg ; Brower:2014daa ; Neuberger:2014pya ; Brower:2020jqj ; Ayyar:2023wub , but recovering rotational invariance of the sphere remained a challenge.
@@
Recently, Ref. Zhu2023 achieved significant progress in this direction via the fuzzy sphere regulator. In a nutshell, they considered a Hamiltonian describing a finite number $N$ of electrons on the surface of $S^{2}$ moving in a constant normal magnetic field and interacting with a short-range potential. Crucially, the model is exactly rotational invariant. Varying the potential, they tuned the system to a quantum phase transition which, in the $N\to\infty$ limit, is in the 3D Ising universality class. The spectrum of the Hamiltonian at criticality was then compared to the 3D Ising CFT, well known due to the conformal bootstrap Simmons-Duffin2017 , showing impressive agreement even for modest values of $N$, and also extracting some energy levels not yet accessed by the bootstrap. Since then, the fuzzy sphere regulator was used to study other Ising CFT observables Hu:2023xak ; Han:2023yyb ; Hu:2023ghk ; Zhou:2023fqu ; Hu:2024pen ; Cuomo:2024psk ; Zhou:2024dbt ; Dedushenko:2024nwi ; Fardelli:2024qla ; Fan:2024vcz and phase transitions in other universality classes Zhou:2023qfi ; Chen:2023xjc ; Han:2023lky ; Chen:2024jxe ; Zhou:2024zud .
@@
While the results of Ref. Zhu2023 represent an excellent baseline for the new fuzzy sphere method, more work is needed to assess how successful this model actually is and to find the best strategy for systematic improvements. This is the first purpose of our work. Here we will be focusing on the original 3D Ising fuzzy sphere model of Zhu2023 , but our considerations are general and can be applied to other models.
@@
We will be relying on Conformal Perturbation Theory (CPT), recently advocated to describe 2+1D models close to but not exactly at their quantum critical point Lao:2023zis ; AMLHStonyBrookTalk . [^2] This framework also proved useful in the recent fuzzy sphere studies Fardelli:2024qla ; Voinea:2024ryq . We will see that CPT provides a comprehensive framework to address many issues arising when comparing a fuzzy sphere model to a CFT, such as: to robustly locate the critical point of the model; to determine the speed of light; to parametrize residual deviation between the microscopic model spectrum and the CFT through an effective theory; or to understand why some models belonging to the same critical line agree with CFT better than others. In particular, we will elucidate what is special about the choice of $V_{0}=4.75$ in Zhu2023 .
@@
With the help of CPT, we will be able to understand and partially subtract finite size corrections, improving the agreement of the fuzzy sphere with the CFT data precisely known thanks to the conformal bootstrap. However, from looking at a larger array of data than what was unveiled in Zhu2023 , we will see that finite size corrections do remain significant for some energy levels.
@@
The second purpose of our work will be to develop a novel method for extracting OPE coefficients of the underlying CFT from the fuzzy sphere model data. Previously, OPE coefficients were extracted Hu:2023xak from a matrix element of a microscopic operator interpolating a CFT operator between eigenstates of the Hamiltonian. Here instead we will note that OPE coefficients of the most relevant $\mathbb{Z}_{2}$ even CFT operator $\varepsilon$ can be extracted by studying the variation of the eigenenergies when the model is detuned from the critical point. We will see that this gives a rather robust scheme. In addition, this allows us to highlight interesting level mixing and repulsion effects which complicate the interpretation of fuzzy sphere data at currently attainable volumes.
@@
Then, in Sec. 4 we test the CPT method in 1+1D. Here we study a 1+1D quantum Hamiltonian with a critical point in the 2D Ising CFT universality class. For a fair comparison, we choose the Hamiltonian not to be exactly solvable. Working on spin chains of length up to $N=26$, we determine the critical point and the speed of light via the CPT method, and we proceed to study residual deviations from the CFT using an effective field theory with three couplings. We find excellent agreement with the CPT predictions derived long ago by Reinicke Reinicke1987a ; Reinicke1987b .
In Sec. 5 we present the numerical results for the fuzzy sphere model in 2+1D and analyze them using the CPT method. Our analysis here differs from Ref. Zhu2023 in several important details, such as:
  • The critical point is determined not through the scaling of the order parameter but by looking at where the CPT coupling $g_{\varepsilon}$ crosses zero Lao:2023zis .
  • While Ref. Zhu2023 privileged the stress tensor level, rescaling the whole spectrum so that the stress tensor scaling dimension is exactly 3, we note that the energy levels $\sigma$ and $\partial\sigma$ are the least affected by perturbations, and fix those levels to their conformal bootstrap values to determine $g_{\varepsilon}$ and the speed of light.
We first study the fuzzy sphere model at the fine-tuned value $V_{0}=4.75$ of the Haldane pseudopotential coefficient found in Ref. Zhu2023 to minimize corrections to scaling. Showing the dependence of energy levels on the system size, we illustrate that in some sectors finite size corrections are significant for attainable system sizes, even for this $V_{0}$. We also study the model at $V_{0}=2.5$ and $V_{0}=6$, to understand better what is special about $V_{0}=4.75$. We found that the two most important irrelevant CPT couplings $g_{\varepsilon^{\prime}}$ and $g_{C}$ are smaller at $V_{0}=4.75$ than away from this value. (The same conclusion for $g_{\varepsilon^{\prime}}$ was reached in Fardelli:2024qla .) In addition, there is no finite-size drift of the critical point at $V_{0}=4.75$, which we hypothesize is related to the vanishing of a curvature contribution to the $g_{\varepsilon}$ coupling (Sec. 5.5.1).
@@
Several detailed discussions are relegated to the appendices. In App. A we provide a guide to the 3D Ising CFT data used in our work. In App. B we describe calculations of relative factors for CPT corrections of descendant states with respect to primary states, and other CPT details. The derivations of all the reported results can be found in the accompanying Mathematica notebook notebook . App. C reviews the 1+1D CPT results of Reinicke Reinicke1987a ; Reinicke1987b . App. D describes the fuzzy sphere model, and App. E the numerical methods used in our work: Exact Diagonalization (ED) for smaller system sizes and Matrix Product States (MPS) for larger ones. We have not used the excellent numerical package FuzzifiED.jl Zhou:2025liv but relied on our own codes.
@@ 2 Ising CFT basics
where $\rho_{0}$ is the microscopic ground-state energy density and $v$ is the speed of light. While not predicted by CFT, they can be determined from a fit. The central charge, scaling dimensions and spins of local CFT operators can then be extracted numerically. This is the essence of the finite-size scaling method for 2D CFTs finite-size-scaling . Note that since momentum is quantized, it receives no corrections: $P_{i}=P^{\rm CFT}_{i}$.
@@
We next review the 3D Ising CFT (Poland:2018epd, , Sec. B.2), which will be the focus of our study using the fuzzy sphere regulator. As in 2D, it is a unitary theory with a global $\mathbb{Z}_{2}$ spin parity and a spatial parity. In 3D there is no Virasoro algebra, as the conformal group in $d\geqslant 3$ is finite dimensional DiFrancesco:1997nk . We have global conformal primaries $\mathcal{O}_{\Delta,\ell}$, where $\Delta\geqslant 0$ is the scaling dimension and $\ell\in\mathbb{Z}_{\geqslant 0}$ is the spin. They are $\ell$-index symmetric traceless tensors, transforming in a $12\ell+$ 1-dimensional representation under SO$(3)$ rotations. The rest of operators organize in global conformal multiplets obtained by acting on primaries with derivatives; these are the descendants.
The 3D Ising CFT is not exactly solved. Only the unit operator $\mathds{1}_{0,0}$ and the stress tensor $T_{3,2}$ have exactly known dimensions; the rest have to be studied numerically. The number of primaries is infinite; the spectrum is dense at high scaling dimensions but sparse at low scaling dimensions. There is also an interesting structure in the spectrum of operators of high spin and low twist $\tau=\Delta-\ell$, which form regular families of “double-twist operators” Simmons-Duffin2017 , whose existence may be understood via analytic continuation in spin Caron-Huot:2017vep .
The 3D Ising CFT spectrum at low scaling dimension, as well as at high spin but low twist, is known with good accuracy thanks to the advances in the numerical conformal bootstrap Rattazzi:2008pe ; El-Showk:2012cjh ; El-Showk:2014dwa ; Kos:2014bka ; Kos:2016ysd ; Simmons-Duffin2017 ; Reehorst:2021hmp ; Chang:2024whx . This includes most importantly the relevant $\mathbb{Z}_{2}$-odd scalar $\sigma$ and the relevant $\mathbb{Z}_{2}$-even scalar $\varepsilon$ of dimensions Chang:2024whx
@@ 2.3 $\mathbb{R}^{d}$/cylinder correspondence
Conformal bootstrap usually studies $d$-dimensional CFTs on $\mathbb{R}^{d}$. Another preferred geometry is the cylinder $S^{d-1}\times\mathbb{R}$. In fact, any CFT quantity can be transferred back and forth between $\mathbb{R}^{d}$ and the cylinder, because these two geometries are related by a Weyl transformation of the metric. [^4] This general $d$ argument takes place of the log map in $d=2$. We will next review this correspondence, which will be very important for our work, following Rychkov:2016iqz ; Simmons-Duffin:2016gjk .
@@
Applying the usual cutting and gluing logic to the path integral Polchinski:1998rq , we can also think of CFT correlators quantum mechanically, both on $\mathbb{R}^{d}$ and on the cylinder (Fig. 2). This point of view will be central for our paper. On $\mathbb{R}^{d}$, one works in the so called “radial quantization,” where the CFT states live on spheres centered at the origin, with the CFT dilatation operator $D$ playing the role of the Hamiltonian. In this picture, inserting the CFT operator $\Phi(0)$ (primary or descendant [^6] We will denote by $\Phi$ a generic CFT operator which may be a primary or a descendant. We will reserve curly letters like ${\cal O}$, ${\cal V}$ etc for primaries.) at the origin prepares the ket state $|\Phi\rangle$ living on the unit sphere, which is an eigenstate of $D$:
@@
These states on the sphere transform under SO$(d)$ rotations as the corresponding CFT operators. In 3D, they therefore form spin-$\ell_{i}$ irreps. On the cylinder, the same CFT states live on the spheres $S^{d-1}$ at constant $\tau$, with $\tau=0$ being the image of the unit sphere in $\mathbb{R}^{d}$. The same ket state at $\tau=0$ can be produced by inserting appropriate cylinder operators $\Phi_{\rm cyl}(\tau=-\infty)$ in the infinite past, although we will not need their exact expressions. [^7] For $\Phi={\cal O}$ primary we have ${\cal O}_{\rm cyl}(\tau=-\infty)=\lim_{\tau\to-\infty}e^{-\tau\Delta_{\cal O}}{\cal O}_{\rm cyl}(\tau,n)$. For $\Phi$ descendant, $\Phi_{\rm cyl}(\tau=-\infty)$ has to be obtained by differentiating (5).
@@
We saw in (1) that $w=-c/12$ in 2D; see Cappelli:1988vw for the 4D case. In 3D as in all odd $d$ there is no Weyl anomaly and we have $w=0$.
@@
where $\Theta^{\rm cyl}$ is a diagonal matrix with $-1$ as the $\tau\tau$ entry, and $1$ for the indices along the sphere, while on $\mathbb{R}^{d}$ we have $\Theta_{\mu\nu}=\delta_{\mu\nu}-x_{\mu}x_{\nu}/x^{2}$.
@@
Now suppose we have a finite microscopic Hamiltonian on the sphere which realizes in the limit of an infinite number $N$ of degrees of freedom a quantum phase transition whose universality class is that of a CFT. $N$ could be a number of spins distributed on the sphere Lao:2023zis . For the fuzzy sphere, $N$ will be a number of electrons. We identify $N$ with the area of the sphere $S^{d-1}_{R}$ such that its radius $R\propto N^{1/(d-1)}$. In such a setup, we expect that the microscopic energy levels on the sphere will be related to the CFT energy levels (12) by the same Eq. (2) as in the 2D case.

### v2

@@ Abstract
Numerical studies of phase transitions in statistical and quantum lattice models provide crucial insights into the corresponding Conformal Field Theories (CFTs). In higher dimensions, comparing finite-volume numerical results to infinite-volume CFT data is facilitated by choosing the sphere $S^{d-1}$ as the spatial manifold. Recently, the fuzzy sphere regulator [1] has enabled such studies with exact rotational invariance, yielding impressive agreement with known 3D Ising CFT predictions, as well as new results. However, systematic improvements and a deeper understanding of finite-size corrections remain essential. In this work, we revisit the fuzzy sphere regulator, focusing on the original Ising model, with two main goals. First, we assess the robustness of this approach using Conformal Perturbation Theory (CPT), to which we provide a detailed guidebook. We demonstrate how CPT provides a unified framework for determining the critical point, the speed of light, and residual deviations from CFT predictions. Applying this framework, we study finite-size corrections and clarify the role of tuning the model in minimizing these effects. Second, we develop a novel method for extracting Operator Product Expansion (OPE) coefficients from fuzzy sphere data. This method leverages the sensitivity of energy levels to detuning from criticality, providing new insights into level mixing and avoided crossings in finite systems. Our work also includes validation of CPT in a 1+1D Ising model away from the integrable limit.
@@ 1 Introduction
Numerical studies of continuous phase transitions in statistical and quantum lattice models are a valuable source of information about the corresponding Conformal Field Theories. In numerical studies done in finite volume, extracting infinite-volume CFT data may be more or less difficult depending on the choice of the spatial manifold. For 1+1D models, studies of quantum Hamiltonians and transfer matrices on a circle $S^{1}$ allow for easy comparison to CFT thanks to the radial quantization [2, 3, 4, 5]. For higher-dimensional models, analogous simplicity would arise for the spatial manifold being the sphere $S^{d-1}$ [6]. This was studied for 3D lattice models [7, 8, 9, 10, 11, 12], but recovering rotational invariance of the sphere remained a challenge.
@@
Recently, Ref. [1] achieved significant progress in this direction via the fuzzy sphere regulator. In a nutshell, they considered a Hamiltonian describing a finite number $N$ of electrons on the surface of $S^{2}$ moving in a constant normal magnetic field and interacting with a short-range potential. Crucially, the model is exactly rotational invariant. Varying the potential, they tuned the system to a quantum phase transition which, in the $N\to\infty$ limit, is in the 3D Ising universality class. The spectrum of the Hamiltonian at criticality was then compared to the 3D Ising CFT, well known due to the conformal bootstrap [13], showing impressive agreement even for modest values of $N$, and also extracting some energy levels not yet accessed by the bootstrap. Since then, the fuzzy sphere regulator was used to study other Ising CFT observables [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and phase transitions in other universality classes [24, 25, 26, 27, 28].
@@
While the results of Ref. [1] represent an excellent baseline for the new fuzzy sphere method, more work is needed to assess how successful this model actually is and to find the best strategy for systematic improvements. This is the first purpose of our work. Here we will be focusing on the original 3D Ising fuzzy sphere model of [1], but our considerations are general and can be applied to other models.
@@
We will be relying on Conformal Perturbation Theory (CPT), recently advocated to describe 2+1D models close to but not exactly at their quantum critical point [29, 30]. [^2] This framework also proved useful in the recent fuzzy sphere studies [22, 31]. We will see that CPT provides a comprehensive framework to address many issues arising when comparing a fuzzy sphere model to a CFT, such as: to robustly locate the critical point of the model; to determine the speed of light; to parametrize residual deviation between the microscopic model spectrum and the CFT through an effective theory; or to understand why some models belonging to the same critical line agree with CFT better than others. In particular, we will elucidate what is special about the choice of $V_{0}=4.75$ in [1].
@@
With the help of CPT, we will be able to understand and partially subtract finite size corrections, improving the agreement of the fuzzy sphere with the CFT data precisely known thanks to the conformal bootstrap. However, from looking at a larger array of data than what was unveiled in [1], we will see that finite size corrections do remain significant for some energy levels.
@@
The second purpose of our work will be to develop a novel method for extracting OPE coefficients of the underlying CFT from the fuzzy sphere model data. Previously, OPE coefficients were extracted [14] from a matrix element of a microscopic operator interpolating a CFT operator between eigenstates of the Hamiltonian. Here instead we will note that OPE coefficients of the most relevant $\mathbb{Z}_{2}$ even CFT operator $\varepsilon$ can be extracted by studying the variation of the eigenenergies when the model is detuned from the critical point. We will see that this gives a rather robust scheme. In addition, this allows us to highlight interesting level mixing and repulsion effects which complicate the interpretation of fuzzy sphere data at currently attainable volumes.
@@
Then, in Sec. 4 we test the CPT method in 1+1D. Here we study a 1+1D quantum Hamiltonian with a critical point in the 2D Ising CFT universality class. For a fair comparison, we choose the Hamiltonian not to be exactly solvable. Working on spin chains of length up to $N=26$, we determine the critical point and the speed of light via the CPT method, and we proceed to study residual deviations from the CFT using an effective field theory with three couplings. We find excellent agreement with the CPT predictions derived long ago by Reinicke [32, 33].
In Sec. 5 we present the numerical results for the fuzzy sphere model in 2+1D and analyze them using the CPT method. Our analysis here differs from Ref. [1] in several important details, such as:
  • The critical point is determined not through the scaling of the order parameter but by looking at where the CPT coupling $g_{\varepsilon}$ crosses zero [29].
  • While Ref. [1] privileged the stress tensor level, rescaling the whole spectrum so that the stress tensor scaling dimension is exactly 3, we note that the energy levels $\sigma$ and $\partial\sigma$ are the least affected by perturbations, and fix those levels to their conformal bootstrap values to determine $g_{\varepsilon}$ and the speed of light.
We first study the fuzzy sphere model at the fine-tuned value $V_{0}=4.75$ of the Haldane pseudopotential coefficient found in Ref. [1] to minimize corrections to scaling. Showing the dependence of energy levels on the system size, we illustrate that in some sectors finite size corrections are significant for attainable system sizes, even for this $V_{0}$. We also study the model at $V_{0}=2.5$ and $V_{0}=6$, to understand better what is special about $V_{0}=4.75$. We found that the two most important irrelevant CPT couplings $g_{\varepsilon^{\prime}}$ and $g_{C}$ are smaller at $V_{0}=4.75$ than away from this value. (The same conclusion for $g_{\varepsilon^{\prime}}$ was reached in [22].) In addition, there is no finite-size drift of the critical point at $V_{0}=4.75$, which we hypothesize is related to the vanishing of a curvature contribution to the $g_{\varepsilon}$ coupling (Sec. 5.5.1).
@@
Several detailed discussions are relegated to the appendices. In App. A we provide a guide to the 3D Ising CFT data used in our work. In App. B we describe calculations of relative factors for CPT corrections of descendant states with respect to primary states, and other CPT details. The derivations of all the reported results can be found in the accompanying Mathematica notebook [34]. App. C reviews the 1+1D CPT results of Reinicke [32, 33]. App. D describes the fuzzy sphere model, and App. E the numerical methods used in our work: Exact Diagonalization (ED) for smaller system sizes and Matrix Product States (MPS) for larger ones. We have not used the excellent numerical package FuzzifiED.jl [35] but relied on our own codes.
@@ 2 Ising CFT basics
where $\rho_{0}$ is the microscopic ground-state energy density and $v$ is the speed of light. While not predicted by CFT, they can be determined from a fit. The central charge, scaling dimensions and spins of local CFT operators can then be extracted numerically. This is the essence of the finite-size scaling method for 2D CFTs [5]. Note that since momentum is quantized, it receives no corrections: $P_{i}=P^{\rm CFT}_{i}$.
@@
[…]
The 3D Ising CFT is not exactly solved. Only the unit operator $\mathds{1}_{0,0}$ and the stress tensor $T_{3,2}$ have exactly known dimensions; the rest have to be studied numerically. The number of primaries is infinite; the spectrum is dense at high scaling dimensions but sparse at low scaling dimensions. There is also an interesting structure in the spectrum of operators of high spin and low twist $\tau=\Delta$-$\ell$, which form regular families of “double-twist operators” [13], whose existence may be understood via analytic continuation in spin [38].
The 3D Ising CFT spectrum at low scaling dimension, as well as at high spin but low twist, is known with good accuracy thanks to the advances in the numerical conformal bootstrap [39, 40, 41, 42, 43, 13, 44, 45]. This includes most importantly the relevant $\mathbb{Z}_{2}$-odd scalar $\sigma$ and the relevant $\mathbb{Z}_{2}$-even scalar $\varepsilon$ of dimensions [45]
@@ 2.3 $\mathbb{R}^{d}$/ cylinder correspondence
Conformal bootstrap usually studies $d$-dimensional CFTs on $\mathbb{R}^{d}$. Another preferred geometry is the cylinder $S^{d-1}\times\mathbb{R}$. In fact, any CFT quantity can be transferred back and forth between $\mathbb{R}^{d}$ and the cylinder, because these two geometries are related by a Weyl transformation of the metric. [^4] This general $d$ argument takes place of the log map in $d=2$. We will next review this correspondence, which will be very important for our work, following [46, 47].
@@
Applying the usual cutting and gluing logic to the path integral [48], we can also think of CFT correlators quantum mechanically, both on $\mathbb{R}^{d}$ and on the cylinder (Fig. 2). This point of view will be central for our paper. On $\mathbb{R}^{d}$, one works in the so called “radial quantization,” where the CFT states live on spheres centered at the origin, with the CFT dilatation operator $D$ playing the role of the Hamiltonian. In this picture, inserting the CFT operator $\Phi(0)$ (primary or descendant [^6] We will denote by $\Phi$ a generic CFT operator which may be a primary or a descendant. We will reserve curly letters like $\cal O$}, $\cal V$} etc for primaries.) at the origin prepares the ket state $|\Phi\rangle$ living on the unit sphere, which is an eigenstate of $D$:
@@
These states on the sphere transform under SO(d)(d) rotations as the corresponding CFT operators. In 3D, they therefore form spin-$\ell_{i}$ irreps. On the cylinder, the same CFT states live on the spheres $S^{d-1}$ at constant $\tau$, with $\tau=0$ being the image of the unit sphere in $\mathbb{R}^{d}$. The same ket state at $\tau=0$ can be produced by inserting appropriate cylinder operators $\Phi_{\rm cyl}(\tau=-\infty)$ in the infinite past, although we will not need their exact expressions. [^7] For $\Phi={\cal O}$ primary we have ${\cal O}_{\rm cyl}(\tau=-\infty)=\lim_{\tau\to-\infty}e^{-\tau\Delta_{\cal O}}{\cal O}_{\rm cyl}(\tau,n)$. For $\Phi$ descendant, $\Phi_{\rm cyl}(\tau=-\infty)$ has to be obtained by differentiating (5).
@@
We saw in (1) that w=c/12w=-c/12 in 2D; see [49] for the 4D case. In 3D as in all odd $d$ there is no Weyl anomaly and we have $w=0$.
@@
where $\Theta^{\rm cyl}$ is a diagonal matrix with 1-1 as the $\tau\tau$ entry, and 11 for the indices along the sphere, while on $\mathbb{R}^{d}$ we have $\Theta_{\mu\nu}=\delta_{\mu\nu}-x_{\mu}x_{\nu}/x^{2}$.
@@
Now suppose we have a finite microscopic Hamiltonian on the sphere which realizes in the limit of an infinite number $N$ of degrees of freedom a quantum phase transition whose universality class is that of a CFT. $N$ could be a number of spins distributed on the sphere [29]. For the fuzzy sphere, $N$ is the number of available magnetic orbitals, not taking into account the spin degeneracy. For the model we consider (see App.D for more details), it also corresponds to the number of electrons. The number of orbitals is directly proportional to the area of the sphere $S^{d-1}_{R}$, such that its radius $R\propto N^{1/(d-1)}$. This corresponds to a physical limit of constant magnetic field on the surface of the sphere when scaling $N$. In such a setup, we expect that the microscopic energy levels on the sphere will be related to the CFT energy levels (12) by the same Eq. (2) as in the 2D case.